# TTQA-RS- A break-down prompting approach for Multi-hop Table-Text Question Answering with Reasoning and Summarization

### v1

@@ 5.2 Analysis and Ablation Studies
This section describes all the analysis and ablation studies performed on our model. Table 3 shows the ablation studies of our model using HybridQA dataset and OTT-QA’s development set. We can observe that baseline CoT model outperforms the baseline standard prompting model. This shows the importance of reasoning in the multi-hop table-text QA task. Then, we test the model by adding the entity type prediction of the expected answer in the CoT prompt. We notice a significant increase in the performance of the model for all the datasets. In the 4th row of Table 3 we have added all the components of our final model except the table-text summary and we can see a further increase in performance in the HybridQA and OTT-QA datasets. Finally, we show the performance of our model TTQA-RS (i.e. last row) by including the generated table-text summary, and we can observe that our model performs the best with all the steps included (including summarization). Adding the table-text summary in the QA input prompt, helps the LLM model to recognize relevant information related to the table or text passages that might have otherwise gone unnoticed. This shows the importance of every component of our model and the need for our break-down prompting approach for table-text QA. In Appendix  C, we evaluated the impact of the number of shots on the model’s performance.
@@ 5.3 Human Evaluation Results
We have manually evaluated the first 100 samples of the table-text summaries generated by LLaMA3-70B, and also the independent sub-questions generated using LLM prompting. We obtained an accuracy of 91% for question decomposition.
Table  4 tabulated the human evaluation results of the generated summaries for the sampled HybridQA test set. For evaluating the generated summaries using retrieved table rows and passages, we have used three evaluation metrics - correctness, inclusivity, and completeness. For correctness, we checked if the summary generated is overall correct and if the model generates any hallucination. For inclusivity, we checked if the generated summaries included information about both the retrieved rows and passages. Completeness was used to check if the generated summaries had complete sentences. We have included all our human evaluation results in the Supplementary section.
@@ 6 Conclusion
This paper proposes a prompting strategy of multi-hop table-text QA by generating table-text summaries and answers of sub-questions. We show that including summaries of retrieved table rows and passages in the prompt with our breakdown approach can substantially increase the performance of CoT prompting in table-text QA. The proposed method achieves new state-of-the-art performance among the prompting approaches for multi-hop table-text QA tasks using both open-source (i.e. LLaMA3-70B) and GPT-4 models. Our experiments specifically focussed on improving prompting strategies in the table-text QA readers. In the future, we plan to extend our work on the table-text QA retrievers which can further improve the QA performance.
@@
Our work has several limitations. Firstly, we are breaking down our problem into individual steps. Even though breaking down the problem into sub-problems helps to reduce hallucination while reasoning with open-source LLMs, it also causes error propagation. Errors made in the initial steps can result in wrong answers. Furthermore, in the experiment of using GPT-4 on our model, we have limited its usage only to the last step of the reader model as GPT-4 is expensive. Using GPT-4 in the remaining steps of the reader could have further improved the performance of our model.
Secondly, the performance of our prompting-based approach, even though is on par with the fine-tuned state-of-the-art models (or has outperformed the training-based state-of-the-model for OTT-QA development set), it’s performance is still not close to the human performance. Also, we are using an existing retriever and the focus of this paper has only been to improve the reader’s performance for multi-hop table-text QA. There is still potential to improve the overall performance of the model by using a better table and text retriever for this problem. Also, currently, we have only experimented with multi-hop table-text datasets in which the questions are already linked to the tables. The test set of the OTT-QA dataset does not have links between the tables with texts. This is out of scope of this current work, but in the future, we plan to explore more in this area.
Acknowledgments
We would like to express our gratitude to Dzmitry Kasinets, a PhD student in the Department of Computer and Information Science and Engineering from University of Florida, for his assistance with the human evaluation of the LLM-generated table-text summaries and independent sub-questions.
@@ Appendix A Implementation Details
For all our experiments we have made use of Nvidia Geforce GTX 1660 Ti. We have set the temperature to 0.5 for all experiments that uses zero-shot or few-shot learning. The table-text summary was generated using zero-shot learning with LLaMA 3-70B and for the question decomposition step, we used two-shot learning with LLAMA3B-70B. This was consistent for all our experiments. In the last two stages of our TTQA-RS framework, i.e. for steps involving table-text QA of independent sub-questions and for table-text QA of original questions, we have used few-shot learning with CoT, and we have experimented with different language models such as LLaMA 2-7b, LLaMA2-13b, LLaMA2-70b, LLaMA3-8b, and LLaMA3-70b.
@@ Appendix B Table-text summarization prompt
Appendix C Impact of number of shots

### v2

@@ 5.2 Analysis and Ablation Studies
5.2.1 Effect of converting rows to sentences
We assessed the retriever model’s effectiveness on the HybridQA test set under two conditions: first, when the table rows are converted into sentences prior to retrieval, and second, when the table rows are used in their original form without conversion. We obtained HIT@1 of 71.49 when rows were converted to sentences and HIT@1 of 60.75 when the tables were directly fed to the table retriever. This shows that converting the rows to sentences before feeding them to the retriever improves the performance of our table-text QA retriever model.
5.2.2 Effect of reader
Table 4 shows the ablation studies of our model using HybridQA dataset and OTT-QA’s development set. We can observe that baseline CoT model outperforms the baseline standard prompting model. We tested the model by adding each component of our reader one by one, and we notice a significant increase in the performance of the model for all the datasets. This shows the importance of every component of our model. In Appendix  D, we evaluated the impact of the number of shots on the model’s performance.
@@ 5.3 Human Evaluation Results
We manually evaluated the first 80 samples of the table-text summaries generated by LLaMA3-70B, and 100 samples of the independent sub-questions generated using LLM prompting were also evaluated. We obtained an accuracy of 91% for question decomposition. For evaluating the generated summaries using retrieved table rows and passages, we have used three evaluation metrics - correctness, inclusivity, and completeness. For correctness, we checked if the summary generated is overall correct and if the model generates any hallucination. For inclusivity, we checked if the generated summaries included information about both the retrieved rows and passages. Completeness was used to check if the generated summaries had complete sentences. We obtained a correctness score, inclusivity score, and completeness score of 0.83, 0.87, and 0.98, respectively, for the 80 HybridQA test set samples.
@@ 6 Conclusion
This paper proposes a RAG-based framework with an enhanced retriever by converting rows to sentences before feeding the tables to a retriever. We have further proposed a prompting strategy of multi-hop table-text QA by generating table-text summaries and answers of sub-questions. We show that including summaries of retrieved table rows and passages in the prompt with our breakdown approach can substantially increase the performance of CoT prompting in table-text QA. The proposed method achieves new state-of-the-art performance among the prompting approaches for multi-hop table-text QA tasks using open-source (i.e. LLaMA3-70B) models.
@@
Our work has several limitations. Firstly, we are breaking down our problem into individual steps. Even though breaking down the problem into sub-problems helps to reduce hallucination while reasoning with open-source LLMs, it also causes error propagation. Errors made in the initial steps can result in wrong answers. Secondly, the performance of our prompting-based approach, even though is on par with the fine-tuned state-of-the-art models (or has outperformed the training-based state-of-the-model for OTT-QA development set), it’s performance is still not close to the supervised models (in case of HybridQA dataset) and also not close to human performance. Also, currently, we have only experimented with multi-hop table-text datasets in which the questions are already linked to the tables. The test set of the OTT-QA dataset does not have links between the tables with texts. This is out of scope of this current work, but in the future, we plan to explore more in this area.
@@ Appendix A Implementation Details
For all our experiments, we used Nvidia Geforce GTX 1660 Ti. For the retriever framework, we used a pre-trained BART-Large model for converting rows to sentences, with a maximum length set to 50, and the number of returned sequences set to 1. For the table retriever, the DPR is fine-tuned with batch size of 32 and trained for 8 epochs, with a learning rate of 2e-5. We trained the model for 5000 questions which took 6 hours to train the model. For the re-ranker model, we have used the pre-trained ’deepset/roberta-base-squad2’ model with the maximum length set to 512, and number of returned sequences set to 1. We have set the $\alpha$ to 0.2 for our experiment.
For the reader, we have set the temperature to 0.5 for all experiments that use zero-shot or few-shot learning. The table-text summary was generated using zero-shot learning with LLaMA 3-70B and for the question decomposition step, we used two-shot learning with LLAMA3B-70B. This was consistent for all our experiments. In the last two stages of our TTQA-RS framework, i.e. for steps involving table-text QA of independent sub-questions and for table-text QA of original questions, we have used few-shot learning with CoT, and we have experimented with different language models such as LLaMA 2-7b, LLaMA2-13b, LLaMA2-70b, LLaMA3-8b, and LLaMA3-70b.
@@ Appendix B Table-text summarization prompt
Appendix C Independent sub-questions
Table  5 shows examples of original complex questions and their corresponding independent sub-questions. The sub-answer obtained using the independent sub-question aids in answering the original complex question.
Table 5: Examples of complex questions and independent sub-questions
Original Complex Question Independent sub-question
What was the release date of the game which Andrew Voss provided commentary on ? What game has Andrew Voss provided commentary on ?
What is the population of the hometown of the 2012 Gatorade Player of the Year ? What is hometown of the 2012 Gatorade Player of the Year ?
Which battle took place in largest country in sub-Saharan Africa ? What is the largest country in sub-Saharan Africa ?
What did the Japanese rename the road that Hop Yat Church is located along ? Where is Hop Yat Church located?
What is the genre of the game that began development as Street Fighter II , but had its name and genre changed ? Which game began development as Street Fighter II, but had its name and genre changed?
What was the highest legal job of the chancellor of Wilfrid Laurier University after May 1990 ? Who was the chancellor of Wilfrid Laurier University after May 1990?
What is the availability status of the Buffalo station owned by the E. W. Scripps Company ? Which station in Buffalo is owned by the E. W. Scripps Company?
Which war film was directed by the director of Hibiscus Town ? Who directed Hibiscus Town?
How many plays has the person known for The Black Coat written ? Who is the person known for The Black Coat?
Appendix D Impact of number of shots